A muscle builder holds the ends of a massless rope. At the center of the rope,

a 15-kg ball is hung as shown. What is the tension in the rope if the angle θ in the drawing is 4.5° from the horizontal?



1
Expert's answer
2021-11-08T09:11:59-0500

"2\cdot T\cdot \sin4.5°=mg\to T=\frac{mg}{2\cdot\sin4.5°}=\frac{15\cdot9.8}{2\cdot\sin4.5°}\approx940\ (N)" . Answer